# Learning large softmax mixtures with warm start EM

### v1

@@ 1.1 Our contribution: learning softmax mixtures with many support points
    EM-based parameter estimation with MoM warm-start, in softmax mixtures.
As a snap-shot of our results, we will see that the MoM estimator’s rates of convergence imply that we can consistently estimate well separated mixture parameters when $p,N\to\infty$. From a theoretical standpoint, these results suggest that MoM yields a good estimator; however, its empirical performance is poor unless $p$ is very large, and the procedure is numerically unstable. This is in line with the previously observed numerical behavior of the MoM estimators in other mixture models, see Lindsay and Basak (1993) for instance. But the MoM procedure does yield estimators in a large neighborhood of the target, even for moderate values of $p$. That makes them excellent warm start candidates for iterative algorithms. We therefore develop an EM algorithm, initialized via MoM. The challenges associated with the analysis of the overall procedure combine those associated with MoM and EM, and we note the source of each difficulty. The construction of the MoM estimators requires new ideas, but their analysis, although also new, is relatively standard. On the other hand, the iterates of the EM algorithm have relatively standard form, but their analysis is surprisingly involved. This work provides a complete analysis of softmax mixture parameter estimation via EM with a MoM warm start. We show that the combination of the two procedures yields a polynomial-time estimator for the parameters of softmax mixtures, which enjoys nearly optimal rates.
A Method of Moments estimator for Softmax Mixture parameters (MoM)
It has been understood for several decades now that the parameters of finite mixtures can be uniquely recovered from moments of the mixing measure
$\rho:=\sum_{k=1}^{K}\alpha_{k}\delta_{{{\bm{\theta}}}_{k}},$
where $\delta_{a}$ denotes the Dirac measure supported on a point $a$. In Lemma 1 of Section 2.1 below we collect existing results, stemming from the seminal works of Lindsay (1989), for univariate mixtures, and of Lindsay and Basak (1993), for multivariate mixtures. Lemma 1 is also the basis of a MoM estimator construction: the parameters of the mixture are given explicitly as solutions of equations involving moments of $\rho$. A method of moments estimator can be therefore considered as long as these moments can be estimated from the observed data, noting that one does not have access to samples from $\rho$.
Estimation of the moments of $\rho$, henceforth referred to as latent moments, from i.i.d. samples from a mixture, is a well understood problem only when the mixture components belong to certain classes of distributions. In the one-dimensional case, we refer to Lindsay (1989) for an overview of the so-called quadratic variance exponential families, with the Gaussian distribution as a chief example. Owing to the properties of these distributions, key results that allow for latent moment estimation from a sample from the mixture distribution have been known for a long time, see for instance Tucker (1963) and Brockett (1977), and have been widely applied, for example, to mixtures of Gaussian (Wu and Yang, 2020) and Binomial (Tian et al., 2017) distributions. Extensions of these ideas to multivariate distributions have been exploited, to the best of our knowledge, only in the context of multivariate Gaussian mixtures. We refer to Lindsay and Basak (1993), for parameter estimation, and to the much more recent result of Doss et al. (2023), who focus on estimation of the mixing measure $\rho$.
In particular, latent moment estimation, from data generated as in (1) and (4), for the softmax mixture given by (2) and (3) have not been studied. In the notation of Lemma 1 below, the task will be to estimate the $r$-th moments $m_{r}$ of the projection of $\rho$ onto one coordinate, and also the mixed moments $m_{ri}$ introduced in Section 2.1. We propose to achieve it in two stages. The first stage is key, and it is one of our main contributions. It consists in the construction of quasi-estimators of these moments, using the sample (1). For instance, a quasi-estimator of $m_{r}$ will be $\bar{m}_{r}$, given by (16) in Section 2.2. We then estimate $\bar{m}_{r}$ by $\widehat{m}_{r}$ given by (14), using the sample (4) from the mixture, thereby completing the estimation of the latent moments $m_{r}$. Once the crucial step of estimating the latent moments is performed, one can appeal to established method of moments parameter estimation strategies. For instance, we use Lemma 1 to obtain estimates of $\bm{\alpha}$ and ${\bm{\theta}}_{k}$, $k\in[K]$, with latent moments replaced by estimated moments and projected moments, as explained in Section 2.3.
Rates of convergence for multivariate mixture parameters are understudied, and the few existing results are restricted to Gaussian mixtures. In Lindsay and Basak (1993), consistency of their MoM parameter estimator is alluded to, but without quantitative rates. The only work specifically mentioning rates for parameter estimation is Doss et al. (2023), but their procedure is developed for their goal, minimax rate optimal mixing measure estimation in the Wasserstein distance, not rate-optimal parameter estimation.
Though the rates we prove for parameter estimation (Propsition 4) show that MoM consistently estimates the true parameters as $p,N\to\infty$, this estimator has several drawbacks. First, implementing MoM requires knowledge of a direction $v\in\mathbb{S}^{L-1}$ such that the projections of the parameters ${\bm{\theta}}_{1}^{*},\dots,{\bm{\theta}}_{K}^{*}$ to the subspace spanned by $v$ are well separated. While it is possible to obtain a weak guarantee by choosing $v$ randomly, the resulting rates of estimation exhibit suboptimal scaling in $L$, see Section 2.5. This motivates augmenting the MoM estimator with a preliminary dimension-reduction step to replace the ambient dimension $L$ with $K$-dimensional subspace spanned by ${\bm{\theta}}_{1}^{*},\dots,{\bm{\theta}}_{K}^{*}$, which improves the dependence of the bound on $L$. Second, even after dimension reduction, our upper bounds on the performance of the MoM estimator include large pre-factors depending on $K$, the norm of ${\bm{\theta}}^{*}_{1},\dots,{\bm{\theta}}^{*}_{K}$, and the separation between the parameters. Therefore, though the MoM estimator is consistent, we observe empirically that its performance can be poor. We correct this by making use of the MoM estimator as a warm start of an EM algorithm. The resulting estimator achieves nearly optimal rates and has much better performance in simulations.
An EM estimator of softmax mixture parameters with MoM initialization
In general, MoM is computationally fast and provides consistent estimates under mild conditions, but it is oftentimes outperformed by the maximum likelihood estimator (MLE). However, in mixture models such as the softmax mixtures that we consider in this paper, the MLE is generally intractable due to the non-concave nature of the log-likelihood function. The EM algorithm (Dempster et al., 1977) is designed to iteratively approximate the log-likelihood for such cases and has been observed to yield satisfactory performance in a wide range of applications. In our problem, the log-likelihood is non-concave in ${\bm{\omega}}=({\bm{\alpha}},{\bm{\theta}}_{1},\ldots,{\bm{\theta}}_{K})$ and we replace it by a convex surrogate function $\widehat{Q}({\bm{\omega}}\mid{\bm{\omega}}^{\prime})$ that is explicitly derived in (40). Evaluating this surrogate function per iteration corresponds to the “E-step”, while maximizing over ${\bm{\omega}}$ is the “M-step”. Since maximization over ${\bm{\alpha}}$ yields a closed-form solution, whereas maximizing over ${\bm{\theta}}_{k}$ does not, we propose in Section 3.1 a hybrid EM algorithm that keeps updating ${\bm{\alpha}}$ by its closed-form solution in (42), while updating ${\bm{\theta}}_{k}$ via the first-order gradient ascent step in (3.1), until convergence.

### v2

@@ 1.1 The softmax mixture model
    We develop a new Method of Moments (MoM), specifically tailored to softmax mixtures, for estimating the latent moments of the mixing measure $\rho:=\sum_{k=1}^{K}\alpha_{k}^{*}\delta_{{\bm{\theta}}_{k}^{*}}$, where $\delta$ denotes the Dirac measure on $\mathbb{R}^{L}$. Under the assumption that the features $x_{1},\ldots,x_{p}$ are independent realizations from a given distribution, we make use of this construction in three related, but different, ways. The background is given in Section 1.2.2.
    • We use a system of equations involving appropriate latent moment approximations, at the population level, to find initial atoms and weights close to the true parameters. Those are then used to initialize a population level EM algorithm to prove that softmax mixtures are globally identifiable, for $p$ large enough. This is the content of Sections 3.1 and 3.2.
    • We develop the sample level analogue of this result. We derive MoM parameter estimates in softmax mixtures, and offer the first rate analysis under this model. The analysis complements that for Gaussian Mixture Models ($K$-GMM), and is valid for any $L\geq 1$. We show, in Section 3.3 that MoM estimators can serve as a warm start for the EM algorithm, but their performance deteriorates fast as $K$ increases.
    • We recommend random initialization when $K$ is not small. For this, in Section 3.4 we develop an estimator for the subspace spanned by ${\bm{\theta}}_{1},\ldots,{\bm{\theta}}_{K}$, tailored to softmax mixtures, and based only on second-order latent moment estimates. We show how to use this subspace estimator to reduce the number of random draws needed to initialize the EM algorithm.
In addition to bridging the existing theoretical and algorithmic gap in softmax mixture estimation via the EM algorithm, our focus on parameter estimation is also motivated by the model’s applications. The model is widely used in the econometrics literature, and could also play an important role in understanding aspects of an LLM output. We give below instances of such applications.
Basic discrete choice models. Softmax mixtures were introduced in the econometrics literature by Boyd and Mellman (1980) and Cardell and Dunbar (1980) under the name “mixed multinomial logits” to model the preference of a heterogeneous set of consumers for a set of mutually exclusive goods. In this application, each vector $x_{j}\in\mathbb{R}^{L}$ reflects the set of attributes of each of the $p$ different goods, while the vector ${{\bm{\theta}}}$ reflects a customer’s preferences for each attribute. The model posits that customers act via random utility maximization: the customer chooses good $j^{\star}=\operatorname*{argmax}_{j\in[p]}x_{j}^{\top}{{\bm{\theta}}}+\epsilon_{j}$. Here $\epsilon_{1},\ldots,\epsilon_{p}$ are independent stochastic terms that reflect idiosyncratic variations in the consumer’s taste. When $\epsilon_{1},\ldots,\epsilon_{p}$ are chosen to have a Gumbel distribution, then (see, e.g., Yellott Jr, 1977)
$\mathbb{P}\left\{j^{\star}=j\right\}=\frac{\exp(x_{j}^{\top}{\bm{\theta}})}{\sum_{i=1}^{p}\exp(x_{i}^{\top}{\bm{\theta}})}\quad\quad\text{for each $j\in[p]$,}$
so that a customer with preference vector $\bm{{\bm{\theta}}}$ chooses among the observed goods $x_{1},\ldots,x_{p}$ according to the probability vector $A({\bm{\theta}})$. This is an appropriate model for the choices of a single customer (or, more generally, for a group of customers with identical preferences). To model the behavior of a large number of consumers with heterogeneous preferences, Boyd and Mellman (1980) and Cardell and Dunbar (1980) suggested to model the population as consisting of a mixture of consumers with different taste vectors. The aggregate probabilities of individual goods being selected is then given by the softmax mixture model (2). This model has been broadly adopted throughout the management science and econometrics literature due to its flexibility and practicality, see (McFadden and Train, 2000; Johnston et al., 2017; Train, 2009; Cameron and Trivedi, 2005) and references therein.
Next word prediction in LLM. Open ended text continuation via LLM is now routinely obtained in response to a prompt of interest, one word at a time. Formally, the prompt is tokenized to yield $u_{1},\ldots,u_{m}\in\mathbb{R}^{L}$, for some initial values of these vectors. This sequence is run through a transformer-based model, initially introduced by Vaswani et al. (2017), to yield contextually embedded vectors $z_{1},\ldots,z_{m}$, of which one is chosen, say $z\in\mathbb{R}^{L}$. Given a vocabulary $x_{1},\ldots,x_{p}$ of vectors in $\mathbb{R}^{L}$ that are viewed as identifiers of the $p$ possible next words (we use tokens and words interchangeably here, although tokens are typically smaller units), the next predicted word is obtained by drawing from a probability on $p$ words with respective masses given by $A(x_{j}|z):=\exp{(z^{\top}x_{j})}/\sum_{i=1}^{p}\exp{(z^{\top}x_{i})}$, $j\in[p]$. This is the reason behind the well-known fact that re-running the LLM with the same prompt can yield different outcomes. In particular, running this process $N$ times, with the same prompt, will yield a sample $Y_{1},\ldots,Y_{N}$, of potentially different words. This sample can thus be viewed as $N$ independent observations on a discrete random variable $Y$, conditionally on the given $z$. Formally, if $Z$ is a latent, the conditional distribution of $Y$ given $Z=z$ is $A(y|z)$, for $y\in\{x_{1},\ldots,x_{p}\}$. If, further, we seek a summary of the complicated LLM process yielding $z$, we can assume that $Z\sim\rho:=\sum_{k=1}^{K}\alpha_{k}\delta_{{\bm{\theta}}_{k}}$, for ${\bm{\theta}}_{1},\ldots,{\bm{\theta}}_{K}$ being the main directions in $\mathbb{R}^{L}$ explored in order to generate $z$, in response to the initial prompt. Then, the marginal distribution of $Y$ is a softmax mixture,
$Y\sim\pi(y):=\sum_{k=1}^{K}\alpha_{k}{\exp({y^{\top}{\bm{\theta}}_{k}})\over\sum_{j=1}^{p}\exp({{x_{j}}^{\top}{\bm{\theta}}_{k}})},\quad\quad y\in\{x_{1},\ldots,x_{p}\}.$
Estimation of the directions ${\bm{\theta}}_{k}$ and of their respective proportions can be thus used in any additional building block that attempts a correction of the LLM output towards a particular direction.
Finally, we note that our bounds on the rates of estimation of ${\bm{\omega}}^{*}$ trivially imply corresponding error bounds for estimation of $\pi({\bm{\omega}}^{*}):=(\pi(x_{1};{\bm{\omega}}^{*}),\ldots,\pi(x_{p};{\bm{\omega}}^{*}))^{\top}$ via the inequality
$\|\pi({\bm{\omega}})-\pi({\bm{\omega}}^{*})\|_{1}~\leq~\|\bm{\alpha}-\bm{\alpha}^{*}\|_{1}+\max_{k\in[K]}\max_{j\in[p]}|x_{j}^{\top}({\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{*})|\,.$ (3)
Rates of estimation for $\pi({\bm{\omega}}^{*})$ can also be obtained more directly via maximum likelihood estimation (MLE), including through the nonparametric MLE approach (Kiefer and Wolfowitz, 1956), which is known to achieve minimax-optimal rates in related settings (Vinayak et al., 2019). Crucially, however, unlike the estimators we propose and analyze below, the direct computation of the MLE is generally intractable due to the non-concave nature of the log-likelihood function, and there is no known computationally efficient algorithm with sharp theoretical guarantees.
1.2 Our contributions
1.2.1 An EM algorithm for softmax mixtures with provable guarantees
[…]